# Low-curvature quantum corrections from unitary evolution of de Sitter space

### v2

@@ I Introduction
In the example of de Sitter spacetime that we will focus on here, the crucial difference between the different slicings is that a closed slicing is global (i.e., covers the entire spacetime) while the flat slicing only covers half of the spacetime; there is a horizon at which the foliation breaks down. This horizon can be reached in finite time along past-directed geodesics, so that the spacetime becomes geodesically incomplete. In the Hamiltonian picture, one would say that past evolution terminates in finite time. Requiring that the quantum theory remains unitary means that the quantum evolution extends beyond this horizon, necessitating strong departures from the classical evolution in a regime where locally nothing special happens and no local observer would notice the presence of a horizon. Hence, it seems one has to give up either general covariance or unitarity (see also [18]).
@@
In section II we summarise the classical formalism of unimodular gravity for FLRW geometries and written in Ashtekar–Barbero variables. We also review the flat and closed slicings of de Sitter spacetime. Section III discusses the Wheeler–DeWitt quantisation of the model, a choice of Gaussian state and derivation of expectation values and standard deviations in the volume. We show our main result, namely strong quantum effects at the classical horizon. Section IV extends the results to LQC, where we encounter some new features such as a maximum value for the cosmological constant and a quantum correction to $\Lambda$ coming from fundamental discreteness. Qualitative features of our solutions are similar to the Wheeler–DeWitt case. We conclude in section V. Appendix A contains some technical details of the LQC calculations.
@@ II Classical dynamics
II.1 Hamiltonian unimodular gravity with FLRW symmetry
@@
II.2 Foliations of de Sitter spacetime
Let us briefly review the different foliations of de Sitter spacetime; this is standard material that can be found in [32], for example. In cosmology it is more common to adopt cosmic time $t$ (choosing $N=1$) and write the FLRW line element in polar spatial coordinates as
${\rm d}s^{2}=-{\rm d}t^{2}+a^{2}(t)\left[\frac{{\rm d}r^{2}}{1-kr^{2}}+r^{2}{\rm d}\Omega^{2}\right]\,.$ (16)
$D$-dimensional de Sitter space is a spacetime of maximal symmetry, so it has additional isometries beyond the ones of general FLRW spacetimes. There are then different representations of de Sitter space as FLRW geometries, corresponding to different foliations by spatial slices of constant curvature. In particular, there is a global slicing by positively curved ($k=1$) slices which covers the entire spacetime, and another slicing by flat ($k=0$) slices which covers only half of the spacetime. Many more details on various aspects of de Sitter spacetime can also be found, e.g., in [33].
$D$-dimensional de Sitter space can be embedded into a flat $D+1$-dimensional Minkowski spacetime, where it is defined as the timelike hyperboloid of points satisfying $X_{\mu}X^{\mu}=\ell^{2}$ with $\mu=0,\ldots,D$ and $\ell^{2}=3/\Lambda$. This is represented for $D=2$ in figure 1(a). The global coordinates
$X_{0}=\ell\sinh\left(\frac{t}{\ell}\right)\,,\qquad X_{i}=\ell\cosh\left(\frac{t}{\ell}\right)z_{i}\,,$ (17)
where $z_{i}z^{i}=1$, cover the entire hyperbola. With these coordinates the metric is described through the FLRW line element (16) with $a(t)=\ell\cosh(t/\ell)$ and $k=1$, and so this is a closed slicing. On the other hand, there is also the more common flat slicing used here, which can be described through (16) when $k=0$ and $a(t)=e^{t/\ell}$. (To compare with the discussion below (15), recall that $a\sim|\nu|^{1/3}$.) This is accomplished through the choice of coordinates
$X_{0}=\ell\sinh\left(\frac{t}{\ell}\right)+\frac{r^{2}}{2\ell}e^{t/\ell}\,,\qquad X_{1}=\ell\cosh\left(\frac{t}{\ell}\right)-\frac{r^{2}}{2\ell}e^{t/\ell}\,\qquad X_{i}=e^{t/\ell}y_{i}\,,$ (18)
where now $i$ ranges from $2$ to $D$, and $y_{i}y^{i}=r^{2}$. Importantly, these coordinates only cover the $X_{0}+X_{1}>0$ half of the hyperbola, as shown in figure 1(b). There is a horizon where constant $t$ slices pile up, and which is reached by past-directed timelike geodesics in finite affine time. In general, no single observer has access to the entire spacetime. The past horizon forms a type of coordinate singularity associated to a particular foliation. As we have already seen (cf. (15)), the horizon can be defined to be at $T=0$ with respect to unimodular time.
Figure 1: (a) 2-dimensional de Sitter visualised as a hyperboloid in 3-dimensional Minkowski spacetime. (b) In red: region covered by the flat slicing.
@@ III Wheeler–DeWitt quantisation
where $\text{Erf}(x)$ denotes the error function.
@@
In contrast to the results presented here, the analysis of [26, Sec. VI] suggests that the expectation value $\langle|\hat{\nu}|\rangle$, and hence the volume of spatial slices, always remains linear in $T$ following the classical solution (15) for positive $T$, so it does eventually go to zero (and then turns negative, which is impossible). Closer inspection of that paper (e.g., Eq. (6.7)) shows that the authors actually calculate the expectation value of $\hat{\nu}$ rather than of $|\hat{\nu}|$, which would be zero at all times due to the assumed symmetry.

### v3

@@ I Introduction
In the example of de Sitter spacetime that we will focus on here, the crucial difference between the different slicings is that a closed slicing is global (it covers the entire spacetime) while the flat slicing only covers half of the spacetime; there is a horizon at which the foliation breaks down (see, e.g., [18], for details on the different foliations of de Sitter spacetime). This horizon can be reached in finite time along past-directed geodesics, so that the spacetime becomes geodesically incomplete. In the Hamiltonian picture, one would say that past evolution terminates in finite time. Requiring that the quantum theory remains unitary means that the quantum evolution extends beyond this horizon, necessitating strong departures from the classical evolution in a regime where locally nothing special happens and no local observer would notice the presence of a horizon. Hence, it seems one has to give up either general covariance or unitarity (see also [19]).
@@
In section II we summarise the classical formalism of unimodular gravity for FLRW geometries and written in Ashtekar–Barbero variables. Section III discusses the Wheeler–DeWitt quantisation of the model, a choice of Gaussian state and derivation of expectation values and standard deviations in the volume. We show our main result, namely strong quantum effects at the classical horizon. Section IV extends the results to LQC, where we encounter some new features such as a maximum value for the cosmological constant and a quantum correction to $\Lambda$ coming from fundamental discreteness. Qualitative features of our solutions are similar to the Wheeler–DeWitt case. We conclude in section V. Appendix A contains some technical details of the LQC calculations.
@@ III Wheeler–DeWitt quantisation
where $\text{erf}(x)=\frac{2}{\sqrt{\pi}}\int_{0}^{x}{\rm d}y\;e^{-y^{2}}$ denotes the error function.
@@
In contrast to the results presented here, the analysis of [27, Sec. VI] suggests that the expectation value $\langle|\hat{\nu}|\rangle$, and hence the volume of spatial slices, always remains linear in $T$ following the classical solution (15) for positive $T$, so it does eventually go to zero (and then turns negative, which is impossible). Closer inspection of that paper (e.g., Eq. (6.7)) shows that the authors actually calculate the expectation value of $\hat{\nu}$ rather than of $|\hat{\nu}|$, which would be zero at all times due to the assumed symmetry.